Identify the following differential equations and hence solve them

i) y' = -(4/x2) - (y/x) + y2

ii) y = xy' + 1 - ln y'


1
Expert's answer
2021-05-31T09:36:48-0400

"Question(1)\\solve\space riccati's\space equation\\
\frac{dy}{dx}=-\frac{4}{x^2}+y^2 -\frac{y}{x}\\
\\
\\
let y=-\frac{\frac{dv}{dx}}{v}, which\space gives \frac{dy}{dx}=-\frac{\frac{d^2v}{dx^2}}{v}+ \frac{(\frac{dv}{dx})^2}{v^2}:\\
-\frac{\frac{d^2v}{dx^2}}{v}+ \frac{(\frac{dv}{dx})^2}{v^2}= \frac{(\frac{dv}{dx})^2}{v^2}+\frac{\frac{dv}{dx}}{vx}-\frac{4}{x^2}\\
subtract\space \frac{(\frac{dv}{dx})^2}{v^2} from\space both \space sides\space and\space multiply\space by\space v\\
-\frac{d^2v}{dx^2}=\frac{\frac{dv}{dx}}{x}-\frac{4v}{x^2}\\
add \space \frac{d^2v}{dx^2}to \space both \space sides:\\
\frac{d^2v}{dx^2}+\frac{\frac{dv}{dx}}{x}-\frac{4v}{x^2}=0\\
multiply \space both\space sides\space by \space x^2\\
x^2\frac{d^2v}{dx^2}+x \frac{dv}{dx}-4v=0\\
assume\space a\space solution \space to\space this \space euler-cauchy \space equation\space will \space be \space proportional \space to \space x^\lambda \space for \space some \space constant \space \lambda.
\\substitute \space v= x^\lambda \space into \space the \space differential\space equation \\
x^2\frac{d^2 x^\lambda}{dx^2}+x \frac{d x^\lambda}{dx}-4 x^\lambda=0\\
substitute \space \frac{d^2 x^\lambda}{dx^2}=\lambda (\lambda-1)x^{lambda-2}and\space \frac{d x^\lambda}{dx}=\lambda x^{\lambda-1}\\
\lambda^2 x^\lambda -4 x^\lambda=0\\
factor \space out x^\lambda:\\
(\lambda^2 -4 )x^\lambda=0\\
assuming\space x \ne0, the \space zeros \space must \space come \space from\space the \space polynomial:\\
(\lambda^2 -4 )=0\\
(\lambda-2)(\lambda+2)=0\\
\lambda=-2\space or\space\lambda=2
\\the \space root\space \lambda =-2 \space \\gives\space v_1=\frac{c_1}{x^2 }\space as\space a\space solotion , where \space c_1 \space is \space an\space arbitrary\space constant.
\\\\the \space root\space \lambda =2 \space \\gives\space v_2=\frac{c_2}{x^2 }\space as\space a\space solution, where \space c_2 \space is \space an\space arbitrary\space constant.\\the \space general\space solution\space is \space the \space sum\space of\space the \space above\space solutions\space\\
v= v_1 +v_2 =\frac{c_1}{x}+c_2x^2\\
substitute \space back\space for\space y=-\frac{\frac{dv}{dx}}{v}\\
y=\frac{2c_1-2c_2x^4}{c_1x+c_2x^5}\\
adjust \space constant\space and\space simplify\\
y=\frac{2}{x}-\frac{4x^3}{x^4+c_1}\\
-----------------------------------------------\\Question(2)\\
solve\space the\space clairaut\space equation\space \\
y=-log(\frac{dy}{dx})+x\frac{dy}{dx}+1\\
differentiate\space both\space side\space with\space respect \space to\space x\\
\frac{dy}{dx}=x\frac{d^2y}{dx^2}+\frac{dy}{dx}-\frac{\frac{d^2y}{dx^2}}{\frac{dy}{dx}}\\
colloect\space in\space terms\space of\space \frac{d^2y}{dx^2}\\
\frac{dy}{dx}=\frac{dy}{dx}+\frac{d^2y}{dx}(x-\frac{1}{\frac{dy}{dx}})\\
subtact\space \frac{dy}{dx} \space from \space both\space sides\\
\frac{d^2y}{dx}(x-\frac{1}{\frac{dy}{dx}})=0\\
solve\space \frac{d^2y}{dx}=0 \space and \space x-\frac{1}{\frac{dy}{dx}}=0 \space separately\\
for \space \frac{d^2y}{dx}=0\\
integrate \space both\space sides\space with\space wrt \space x\\
\frac{dy}{dx}=\int 0\space dx= c_1 , where \space c_1 is\space an\space arbitrary\space constant.\\
subsitute \frac{dy}{dx}=c_1 into\space y=-log\frac{dy}{dx}+x\frac{dy}{dx}+1\\
y=-log(c_1)+c_1 x+1\\
for \space x-\frac{1}{\frac{dy}{dx}}=0\\
solve \space for\space \frac{dy}{dx}\\
\frac{dy}{dx}=\frac{1}{x}
\\
substitute \space into\space y=-log\frac{dy}{dx}+x\frac{dy}{dx}+1\\
y=-log(\frac{1}{x})+2\\

y=-log(\frac{1}{x})+2 \space or y=-log(c_1)+c_1x+1"

------------------------------------------------------